OpenStudy (anonymous):

when the smallest of 3 consecutive integers is subtracted from twice the largest, the result is 15. What is the middle number?

OpenStudy (anonymous):

12?

OpenStudy (ajprincess):

Let the smallest number be n. then other numbers will be n+1, n+2. Here largest number is n+2. 2(n+2)-n=15 2*n+2*2-n=15 2n+4-n=15 Find n. Then to find the middle number plug in the value of n in n+1.

OpenStudy (anonymous):

i see so that's how you solve it :D mine was not so sure i just guessed afterall

OpenStudy (ajprincess):

@JosellyJs did u find it?

OpenStudy (anonymous):

Thank you so much, yes :)

OpenStudy (ajprincess):

welcome:)